Using Maxwell’s equations in free space, derive the wave equation for the z-component of
the electric field vector.
1
Expert's answer
2019-09-03T12:22:56-0400

The Maxwell's equations in free space


"\nabla\times {\bf E}=-\frac{1}{c}\frac{\partial {\bf B}}{\partial t}""\nabla\times {\bf B}=\frac{1}{c}\frac{\partial {\bf E}}{\partial t}""\rm div{\bf E}=0""\rm div{\bf B}=0"

From the second equation we obtain


"\nabla\times \frac{\partial {\bf B}}{\partial t}=\frac{1}{c}\frac{\partial^2 {\bf E}}{\partial t^2}"

Plug first equation into last, we get


"-c\nabla\times (\nabla\times {\bf E})=\frac{1}{c}\frac{\partial^2 {\bf E}}{\partial t^2}"

Since

"\nabla\times (\nabla\times {\bf E})={\rm grad(div {\bf E})}-\nabla^2 {\bf E} =-\nabla^2 {\bf E}"

we finally obtain


"\nabla^2 {\bf E}-\frac{1}{c^2}\frac{\partial^2 {\bf E}}{\partial t^2}=0"

For the z-component of the electric field vector the wave equation


"\nabla^2 {E_z}-\frac{1}{c^2}\frac{\partial^2 {E_z}}{\partial t^2}=0"